OpenStudy (anonymous):

a1=10 ak+1=(.7)ak what is the value for: S(1) S(10) S(infinity)

OpenStudy (a_clan):

arithmetic progression

OpenStudy (anonymous):

\begin{array}l\color{red}{\normalsize\text{Q}}\color{orange}{\normalsize\text{u}}\color{#9c9a2e}{\normalsize\text{e}}\color{green}{\normalsize\text{s}}\color{blue}{\normalsize\text{t}}\color{purple}{\normalsize\text{i}}\color{purple}{\normalsize\text{o}}\color{red}{\normalsize\text{n}}\color{orange}{\normalsize\text{ }}\color{#9c9a2e}{\normalsize\text{l}}\color{green}{\normalsize\text{o}}\color{blue}{\normalsize\text{o}}\color{purple}{\normalsize\text{k}}\color{purple}{\normalsize\text{s}}\color{red}{\normalsize\text{ }}\color{orange}{\normalsize\text{r}}\color{#9c9a2e}{\normalsize\text{e}}\color{green}{\normalsize\text{a}}\color{blue}{\normalsize\text{l}}\color{purple}{\normalsize\text{l}}\color{purple}{\normalsize\text{y}}\color{red}{\normalsize\text{ }}\color{orange}{\normalsize\text{u}}\color{#9c9a2e}{\normalsize\text{n}}\color{green}{\normalsize\text{f}}\color{blue}{\normalsize\text{a}}\color{purple}{\normalsize\text{m}}\color{purple}{\normalsize\text{i}}\color{red}{\normalsize\text{l}}\color{orange}{\normalsize\text{i}}\color{#9c9a2e}{\normalsize\text{a}}\color{green}{\normalsize\text{r}}\color{blue}{\normalsize\text{}}\end{array}

OpenStudy (a_clan):

first term=10

OpenStudy (anonymous):

Ohhhhhhhhh

OpenStudy (a_clan):

(k+1) th term = 0.7 (kth term)

OpenStudy (anonymous):

I'd have to read up on it

OpenStudy (anonymous):

What happened to the chat box?

OpenStudy (a_clan):

Actually Geometric progression

OpenStudy (anonymous):

Somebody must've baptized it

OpenStudy (a_clan):

\[a _{k+1}/ a _{k} = 0.7 = common ratio\]

OpenStudy (anonymous):

wait a second, did something similar to this in physics with finding ratios and such

OpenStudy (anonymous):

so what is the variable here? a ? we replace a with 1, 10, and infinity?

OpenStudy (a_clan):

first term=a1=10 common ratio = r = a(k+1)/a(k) = 0.7 Sum of n terms, \[\S _{n} = a[1-r ^{n}]/ [1-r]\]

OpenStudy (a_clan):

S(1) = 10 = sum of first 1 terms S(10)= first 10 terms = 10[1-(0.7)^10]/[1-0.7] S(infinity) = a/[1-r] = 10/0.3 = 33.33

OpenStudy (anonymous):

Interesting....I'll have to read more about it.